# SimBEV: A Synthetic Multi-Task Multi-Sensor Driving Data Generation Tool and Dataset

### v1

@@ B.2 SimBEV dataset statistics
Finally, logarithmic BEV ground truth label heat maps of the SimBEV dataset for all classes are shown in Fig. 14. As expected, there is a concentration of road labels in the ego vehicle’s direction of travel, which also results in a concentration of labels in that region for all vehicular classes. In contrast, pedestrian labels are relatively evenly distributed.
C BEVFusion Implementation Details
All variants of BEVFusion [35] were trained on an Nvidia DGX A100 640GB node using sigmoid focal loss [29] and a batch size of 64. We used the original BEVFusion implementation’s settings to train the 3D object detection models, i.e. we employed the AdamW optimizer [38] with a weight decay factor of 0.01. The camera-only, lidar-only, and fused models used an initial learning rate of 2e-4, 1e-4, and 2e-4, respectively. A cyclic learning rate scheduler [52] was used for the first two models, and cosine annealing [37] was used for the latter. All three models were trained for 20 epochs. BEVFusion’s BEV segmentation variants were trained for 40 epochs using an initial learning rate of 3e-5 and a cyclic learning rate scheduler with a 4.0 target ratio. Input data was augmented (translated, rotated, scaled) for the training of all BEVFusion variants.
D Comprehensive Evaluation Results
Class 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Road 64.1 69.1 71.6 73.3 74.1 73.8 72.6 70.5 66.4
Car 6.3 12.7 22.7 28.4 26.7 19.5 15.2 11.6 9.5
Truck 5.7 13.7 18.8 19.7 17.0 12.3 7.2 2.9 0.3
Bus 11.4 23.1 29.1 30.8 29.5 25.6 19.4 11.8 4.5
Motorcycle 0.4 2.5 1.2 0.2 0.0 0.0 0.0 0.0 0.0
Bicycle 0.1 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Rider 0.3 1.0 0.2 0.1 0.0 0.0 0.0 0.0 0.0
Pedestrian 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Mean 11.0 15.3 18.0 19.1 18.4 16.4 14.3 12.1 10.1
Road 52.9 58.6 68.7 87.5 89.6 89.3 87.6 84.6 79.7
Car 6.5 41.0 60.8 72.2 75.8 68.6 57.2 45.0 28.8
Truck 14.7 49.1 66.0 75.1 77.5 71.1 60.2 47.5 29.2
Bus 27.0 65.6 78.8 84.2 84.8 79.5 70.6 60.3 46.2
Motorcycle 6.0 17.4 27.9 38.6 44.6 33.9 9.1 0.3 0.0
Bicycle 2.9 7.8 13.4 12.7 4.0 0.1 0.0 0.0 0.0
Rider 5.2 14.4 24.0 36.2 38.5 13.4 0.3 0.0 0.0
Pedestrian 3.0 11.0 19.7 31.4 27.7 2.1 0.0 0.0 0.0
Mean 14.8 33.1 44.9 54.7 55.3 44.8 35.6 29.7 23.0
Road 75.9 82.9 86.7 89.0 89.9 89.4 87.8 85.2 81.1
Car 20.3 46.1 63.8 73.8 77.0 70.0 59.2 47.5 32.3
Truck 26.1 53.8 66.7 75.5 79.3 73.8 63.2 51.3 33.8
Bus 39.6 69.0 79.9 84.5 84.8 80.3 72.5 63.1 49.7
Motorcycle 5.8 18.8 29.6 39.7 45.7 34.8 9.4 0.1 0.0
Bicycle 3.3 8.7 14.6 17.9 4.5 0.1 0.0 0.0 0.0
Rider 5.4 15.2 24.6 36.2 37.7 15.4 0.9 0.0 0.0
Pedestrian 2.8 11.3 19.9 31.6 32.7 5.8 0.1 0.0 0.0
Mean 22.4 38.2 48.2 56.0 56.5 46.2 36.6 30.9 24.6
Table 11: BEV segmentation IoU values (in %) by class and IoU threshold for the camera-only variant (top), lidar-only variant (middle), and fused variant (bottom) of BEVFusion.
Class 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Car 33.0 29.4 23.8 17.8 12.3 7.1 2.0 0.2 0.0
Truck 30.3 27.0 23.0 18.5 13.8 9.7 5.1 1.5 0.0
Bus 31.5 30.3 27.5 25.1 21.5 15.5 10.0 5.7 0.0
Motorcycle 28.5 25.7 22.0 17.0 11.7 6.4 1.0 0.1 0.0
Bicycle 23.2 19.7 16.6 11.6 6.5 2.6 0.2 0.0 0.0
Pedestrian 2.2 1.0 0.5 0.2 0.1 0.0 0.0 0.0 0.0
Mean 24.8 22.2 18.9 15.0 11.0 6.9 3.1 1.3 0.0
Car 51.4 51.2 50.7 50.3 49.7 48.7 46.1 36.1 5.1
Truck 46.0 45.9 45.7 45.6 45.3 44.7 42.6 30.9 7.3
Bus 35.6 35.4 35.4 35.2 35.0 34.4 34.3 32.1 6.7
Motorcycle 48.1 47.3 46.6 46.0 45.3 44.3 41.5 26.3 1.4
Bicycle 56.5 56.5 56.2 55.8 55.1 52.2 42.3 23.9 0.9
Pedestrian 57.4 56.5 54.8 51.9 46.5 36.7 19.6 3.9 0.1
Mean 49.2 48.8 48.2 47.4 46.1 43.5 37.7 25.5 3.6
Car 51.5 51.3 50.8 50.4 49.8 48.9 46.5 37.2 7.2
Truck 46.2 46.0 45.8 45.7 45.3 44.7 43.4 34.0 11.4
Bus 35.6 35.6 35.5 35.4 35.2 34.9 34.3 32.1 11.8
Motorcycle 48.5 47.9 47.2 46.5 45.8 44.9 42.6 30.3 2.3
Bicycle 56.4 56.3 56.1 55.7 54.9 52.3 43.4 27.1 1.6
Pedestrian 57.3 56.6 54.7 52.0 46.7 37.7 20.8 4.5 0.1
Mean 49.2 48.9 48.4 47.6 46.3 43.9 38.5 27.5 5.7
Table 12: 3D object detection AP values (in %) by class and IoU threshold for the first method (3D IoU-based) for the camera-only variant (top), lidar-only variant (middle), and fused variant (bottom) of BEVFusion.
Class 0.5 1.0 2.0 4.0
Car 10.5 20.3 30.5 37.3
Truck 8.4 15.4 25.4 33.7
Bus 7.0 15.8 24.3 31.6
Motorcycle 16.5 26.4 31.9 34.7
Bicycle 15.6 22.1 27.7 29.7
Pedestrian 4.7 14.2 26.4 35.3
Mean 10.5 19.1 27.7 33.7
Car 34.9 45.0 51.2 51.8
Truck 21.8 36.2 45.9 46.4
Bus 34.3 34.8 35.4 35.6
Motorcycle 48.2 49.5 50.6 50.9
Bicycle 56.4 56.5 56.5 56.6
Pedestrian 38.1 52.8 59.8 60.7
Mean 38.9 45.8 49.9 50.3
Car 34.7 44.9 51.2 51.8
Truck 23.2 36.6 46.3 46.8
Bus 34.0 35.0 35.6 35.6
Motorcycle 48.0 49.1 50.2 50.8
Bicycle 56.3 56.4 56.4 56.4
Pedestrian 38.5 53.0 59.9 60.9
Mean 39.1 45.8 49.9 50.4
Table 13: 3D object detection AP values (in %) by class and distance threshold for the second method (distance-based) for the camera-only variant (top), lidar-only variant (middle), and fused variant (bottom) of BEVFusion.
BEV segmentation IoU values (in %) by class and IoU threshold for the three BEVFusion variants are shown in Tab. 11, and 3D object detection AP values (in %) by class and threshold are shown in Tab. 12 and Tab. 13 for the first and second methods, respectively. As discussed before, the biggest takeaway from the results is that the camera-only variants (for both BEV segmentation and 3D object detection) perform worse than the lidar-only and fused variants. Another takeaway is that compared to the first method, the second method for evaluating 3D object detection results allows for better comparison of the camera-only model results to that of the other two models, especially when it comes to smaller objects.

### v2

@@ B.2 SimBEV dataset statistics
Finally, logarithmic BEV ground truth heat maps for all classes of the SimBEV dataset are shown in Fig. 17. As expected, road is concentrated in the direction of travel of the ego vehicle, which also results in the concentration of labels of all vehicular classes in that region. In contrast, pedestrian labels are relatively evenly distributed.
C 3D Object Detection Evaluation
For both approaches to evaluating the results of 3D object detection, AP is calculated from the area under the precision-recall curve. For the first method, we use IoU thresholds of $\mathbb{T}=\{0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$ to match the bounding boxes. For the second method, similar to [3], we use distance thresholds of $\mathbb{T}=\{0.5,1,2,4\}$ m to match the bounding boxes. For both approaches, we define mAP as the average over all classes and all matching thresholds:
Similar to [3], we measure a set of True Positive metrics (TP metrics) for each predicted bounding box that is matched to a ground truth bounding box: Average Translation Error (ATE), which is the Euclidean distance (in m) between box centers; Average Orientation Error (AOE), which is the smallest yaw angle difference (in rad) between the two boxes; Average Scale Error (ASE), which is equal to one minus the 3D IoU value of the two boxes after aligning for orientation and translation; and Average Velocity Error (AVE), which is the L2 norm of the difference in box velocities (in m/s). The mean TP metric (mTP) for each metric is computed by averaging over all classes and thresholds:
D Model Implementation
All variants of BEVFusion [33] and UniTR [53] were trained on an Nvidia DGX A100 640GB node using the settings and hyperparameters used by their authors for benchmarking on the nuScenes dataset [3]. SimBEV dataset data were augmented (translated, rotated, scaled) during training for all models.
E Comprehensive Evaluation Results
Class Model Modality 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Road BEVFusion-C C 59.5 67.1 71.5 74.5 76.0 75.2 72.6 68.9 62.3
BEVFusion-L L 48.6 55.9 66.1 85.1 87.7 87.2 84.9 81.1 74.6
BEVFusion C + L 59.7 72.0 80.0 85.5 88.4 88.1 85.9 82.4 76.3
UniTR C + L 85.7 89.1 91.0 92.2 92.8 92.5 91.4 89.3 85.5
UniTR+LSS C + L 86.0 89.4 91.3 92.6 93.3 93.0 92.0 90.1 86.4
Car BEVFusion-C C 3.5 8.0 18.8 22.4 17.2 11.3 9.7 8.7 6.3
BEVFusion-L L 5.3 37.8 56.5 67.1 70.6 63.6 51.8 37.6 18.8
BEVFusion C + L 11.7 39.4 58.6 69.4 72.7 65.5 54.0 40.1 20.5
UniTR C + L 31.2 49.6 63.1 71.3 73.8 67.4 57.4 45.8 29.8
UniTR+LSS C + L 32.2 50.6 63.0 70.9 72.8 66.1 55.8 44.3 28.8
Truck BEVFusion-C C 2.1 6.7 11.7 9.8 5.1 2.1 0.4 0.0 0.0
BEVFusion-L L 11.4 44.7 61.2 70.6 73.5 67.4 55.2 39.5 16.3
BEVFusion C + L 12.3 47.0 61.4 70.9 74.5 69.2 57.6 43.2 20.6
UniTR C + L 33.3 51.2 61.7 67.2 67.7 61.3 51.9 40.1 22.7
UniTR+LSS C + L 34.4 53.2 63.6 69.0 69.4 63.4 53.6 41.6 23.6
Bus BEVFusion-C C 2.1 9.0 19.9 24.6 22.9 16.8 10.3 6.0 1.1
BEVFusion-L L 19.1 56.9 72.0 79.7 81.5 78.1 69.7 59.4 44.1
BEVFusion C + L 19.7 56.8 70.3 78.2 80.8 77.2 68.9 59.0 44.1
UniTR C + L 39.5 53.5 56.7 55.5 51.7 45.1 37.8 29.8 17.9
UniTR+LSS C + L 44.4 57.2 61.7 62.0 58.5 51.4 42.9 33.5 21.3
Motorcycle BEVFusion-C C 0.3 0.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BEVFusion-L L 4.8 13.7 23.6 32.7 32.5 15.8 0.8 0.0 0.0
BEVFusion C + L 5.0 13.5 23.6 34.6 36.3 18.3 1.5 0.0 0.0
UniTR C + L 6.2 17.6 29.1 37.4 36.5 22.7 7.2 0.3 0.0
UniTR+LSS C + L 8.3 19.5 29.1 36.3 35.9 21.6 5.8 0.3 0.0
Bicycle BEVFusion-C C 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BEVFusion-L L 1.8 5.1 10.0 13.3 3.6 0.0 0.0 0.0 0.0
BEVFusion C + L 1.9 5.6 11.1 12.7 3.6 0.0 0.0 0.0 0.0
UniTR C + L 3.3 7.9 11.0 13.6 11.4 5.5 0.7 0.0 0.0
UniTR+LSS C + L 3.2 7.7 10.5 10.9 6.3 2.7 1.4 0.2 0.0
Rider BEVFusion-C C 0.3 0.1 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BEVFusion-L L 4.6 11.7 20.8 30.3 18.4 0.5 0.0 0.0 0.0
BEVFusion C + L 4.8 11.9 21.0 31.0 23.3 1.2 0.0 0.0 0.0
UniTR C + L 5.3 15.5 25.7 35.7 36.2 17.4 1.7 0.0 0.0
UniTR+LSS C + L 6.5 15.7 24.4 32.1 31.6 14.8 1.3 0.0 0.0
Pedestrian BEVFusion-C C 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BEVFusion-L L 3.1 9.6 17.6 28.4 18.9 0.1 0.0 0.0 0.0
BEVFusion C + L 3.1 9.9 18.3 28.7 20.2 0.2 0.0 0.0 0.0
UniTR C + L 3.0 11.1 19.9 30.2 27.5 3.9 0.0 0.0 0.0
UniTR+LSS C + L 3.1 10.2 17.7 25.1 12.9 0.2 0.0 0.0 0.0
Mean BEVFusion-C C 8.5 11.4 15.2 16.4 15.2 13.2 11.6 10.5 8.7
BEVFusion-L L 12.4 29.4 41.0 50.9 48.3 39.1 32.8 27.2 19.2
BEVFusion C + L 14.8 32.0 43.0 51.3 50.0 40.0 33.5 28.1 20.2
UniTR C + L 25.9 36.9 44.8 50.4 49.7 39.5 31.0 25.7 19.5
UniTR+LSS C + L 27.3 38.0 45.2 49.9 47.6 39.1 31.6 26.2 20.0
Table 11: BEV segmentation IoUs (in %) by class and IoU threshold for different models evaluated on the SimBEV dataset test set. The top three values are indicated in red, green, and blue, respectively.
Class Model Modality mAP mATE mAOE mASE mAVE
(%) $\uparrow$ (m) $\downarrow$ (rad) $\downarrow$ $\downarrow$ (m/s) $\downarrow$
Car BEVFusion-C C 12.5 0.518 0.710 0.177 5.67
BEVFusion-L L 41.0 0.129 0.080 0.113 1.40
BEVFusion C + L 41.1 0.128 0.078 0.112 1.37
UniTR C + L 38.8 0.100 0.123 0.084 0.49
UniTR+LSS C + L 39.8 0.099 0.106 0.087 0.47
Truck BEVFusion-C C 14.1 0.568 0.902 0.123 6.64
BEVFusion-L L 38.7 0.143 0.040 0.100 1.73
BEVFusion C + L 38.7 0.149 0.042 0.096 1.70
UniTR C + L 36.2 0.108 0.098 0.066 0.56
UniTR+LSS C + L 36.6 0.110 0.096 0.070 0.55
Bus BEVFusion-C C 17.4 0.967 1.225 0.020 5.80
BEVFusion-L L 30.6 0.159 0.044 0.067 2.32
BEVFusion C + L 30.5 0.164 0.037 0.058 2.42
UniTR C + L 25.6 0.114 0.143 0.038 0.85
UniTR+LSS C + L 27.4 0.124 0.121 0.036 0.75
Motorcycle BEVFusion-C C 11.6 0.261 0.688 0.135 6.63
BEVFusion-L L 39.6 0.091 0.080 0.131 1.74
BEVFusion C + L 40.1 0.092 0.071 0.125 1.59
UniTR C + L 39.8 0.074 0.102 0.092 0.54
UniTR+LSS C + L 40.4 0.074 0.097 0.088 0.56
Bicycle BEVFusion-C C 8.4 0.227 0.818 0.200 3.20
BEVFusion-L L 38.4 0.088 0.071 0.186 1.46
BEVFusion C + L 38.3 0.089 0.054 0.177 1.41
UniTR C + L 39.5 0.071 0.110 0.120 0.43
UniTR+LSS C + L 41.3 0.069 0.103 0.103 0.44
Pedestrian BEVFusion-C C 0.2 0.111 1.42 0.035 1.18
BEVFusion-L L 37.0 0.064 0.262 0.076 0.39
BEVFusion C + L 37.2 0.066 0.235 0.069 0.38
UniTR C + L 40.1 0.056 0.376 0.048 0.21
UniTR+LSS C + L 40.5 0.055 0.360 0.049 0.21
Mean BEVFusion-C C 7.0 0.337 0.943 0.106 4.98
BEVFusion-L L 33.9 0.105 0.086 0.107 1.49
BEVFusion C + L 34.1 0.107 0.077 0.101 1.46
UniTR C + L 33.0 0.081 0.140 0.071 0.51
UniTR+LSS C + L 34.2 0.083 0.131 0.069 0.49
Table 12: 3D object detection results for different models evaluated on the SimBEV test set using the first method. The top three values are indicated in red, green, and blue, respectively.
Class Model Modality mAP mATE mAOE mASE mAVE
(%) $\uparrow$ (m) $\downarrow$ (rad) $\downarrow$ $\downarrow$ (m/s) $\downarrow$
Car BEVFusion-C C 23.3 0.824 0.896 0.217 4.94
BEVFusion-L L 46.1 0.165 0.109 0.127 1.44
BEVFusion C + L 46.5 0.162 0.106 0.125 1.40
UniTR C + L 46.5 0.132 0.179 0.095 0.52
UniTR+LSS C + L 46.0 0.128 0.153 0.097 0.50
Truck BEVFusion-C C 20.4 0.751 0.695 0.148 5.55
BEVFusion-L L 46.3 0.162 0.045 0.110 1.75
BEVFusion C + L 46.2 0.168 0.049 0.106 1.74
UniTR C + L 45.2 0.123 0.134 0.074 0.58
UniTR+LSS C + L 45.8 0.128 0.140 0.078 0.56
Bus BEVFusion-C C 18.7 0.829 1.185 0.022 5.55
BEVFusion-L L 34.1 0.169 0.049 0.072 2.40
BEVFusion C + L 34.3 0.176 0.040 0.063 2.44
UniTR C + L 35.1 0.122 0.216 0.041 0.80
UniTR+LSS C + L 35.2 0.129 0.176 0.037 0.72
Motorcycle BEVFusion-C C 26.5 0.604 0.841 0.140 6.63
BEVFusion-L L 51.9 0.114 0.118 0.159 1.79
BEVFusion C + L 51.6 0.113 0.104 0.153 1.65
UniTR C + L 52.3 0.075 0.216 0.105 0.62
UniTR+LSS C + L 52.1 0.091 0.148 0.096 0.61
Bicycle BEVFusion-C C 25.1 0.574 1.117 0.219 4.12
BEVFusion-L L 55.5 0.114 0.087 0.213 1.53
BEVFusion C + L 55.3 0.115 0.073 0.207 1.52
UniTR C + L 54.4 0.087 0.168 0.146 0.49
UniTR+LSS C + L 55.0 0.085 0.155 0.121 0.49
Pedestrian BEVFusion-C C 18.9 0.884 1.529 0.073 1.10
BEVFusion-L L 54.5 0.141 0.392 0.120 0.47
BEVFusion C + L 54.8 0.142 0.362 0.109 0.47
UniTR C + L 52.8 0.118 0.499 0.079 0.29
UniTR+LSS C + L 53.1 0.116 0.472 0.081 0.28
Mean BEVFusion-C C 22.1 0.744 1.044 0.137 4.65
BEVFusion-L L 48.1 0.144 0.133 0.134 1.56
BEVFusion C + L 48.1 0.146 0.122 0.127 1.54
UniTR C + L 47.7 0.113 0.224 0.090 0.55
UniTR+LSS C + L 47.8 0.113 0.207 0.085 0.53
Table 13: 3D object detection results for different models evaluated on the SimBEV test set using the second method. The top three values are indicated in red, green, and blue, respectively.
BEV segmentation IoUs (in %) by class and IoU threshold for different models are shown in Tab. 11, and a breakdown of 3D object detection results by class is shown in Tab. 12 and Tab. 13 for the first and second methods, respectively. As discussed before, the biggest takeaway from the results is that camera-only models (for both BEV segmentation and 3D object detection) perform worse than lidar-only and fusion models.